# Rethinking Score Distillation as a Bridge Between Image Distributions

### v1

@@ Appendix B More Visual Results
Figure 7: 3D sketch-to-real. We introduce a conditional generation task in 3D where a coarse human-drawn mesh is optimized into a high-quality mesh. While SDS and our gradient both adhere to the prompt and shape conditions, our method produces higher fidelity colors and texture.
@@
We first fit an implicit SDF volume to multi-view renders of the mesh, then apply our gradient with the same schedule as in text-based NeRF optimization. We lower the learning rate for geometry parameters to prevent divergence from the guiding sketch. Holding other hyperparameters equal, we compare our gradient and the SDS gradient in Figure 7.
@@
Figure 8: Diffusion illusions. We generate overlaid optic illusions with SDS and our method. While SDS suffers from color artifacts, our methods produce more details and proper color.
@@
Prior works have shown that diffusion models can be leveraged to generate optical illusions [17, 5]. In these settings, the same image looks semantically different when transformed. To use the diffusion model sampling process, a previous study shows that the transformation has to be orthogonal [17]. However, there remain interesting illusions that are not formed by orthogonal transformation. One such is the rotation overlays. Given a base and a rotator image, by composing the base image with the rotator image at different angles, rotation overlays use two images to display four images. As such composition is not defined by an orthogonal matrix, the existing method [5] employs SDS to optimize the base and rotator images. Such a method suffers from the over-saturation problem, as shown in Figure 8. We show that our method can generate such optical illusions with better visual quality.
@@
For text-guided NeRF optimization comparison against baselines, we follow show more results in Fig. 10. We test on the prompts used in the original paper [68] and additional prompts [69] that we find to be challenging. We notice that SDS often suffers from over-saturation problems. Our method does not require training a LoRA while it can still improve SDS by getting rid of the color artifacts and generating more details.
Figure 9: Ablation study of our method without stage 1. We show directly optimizing with $y_{\text{src}}$ from the start could undermine the quality of the geometry and produce unnecessary content.
@@
Instead of switching to stage 2 during the optimization process, we ablate with starting without any SDS optimization from the beginning. That is, we always use the $y_{\text{src}}$ with the descriptors “, oversaturated, smooth, pixelated, cartoon, foggy, hazy, blurry, bad structure, noisy, malformed”. As shown in Figure 9, this makes it hard to generate the proper geometry even though the local texture looks reasonable and is inclined to produce excessive details that are not described by the texts. We suspect that this is because using $y_{\text{src}}$ increases the mismatching error at the beginning of the optimization process when the initialization does not resemble the target prompt at all.
Figure 10: Additional comparison of text-guided NeRF optimization. We show more examples to compare with different distillation methods, SDS and VSD.

### v2

@@ Appendix B More Visual Results
Figure A1: 3D sketch-to-real. We introduce a conditional generation task in 3D where a coarse human-drawn mesh is optimized into a high-quality mesh. While SDS and our gradient both adhere to the prompt and shape conditions, our method produces higher fidelity colors and texture.
B.1 Additional Applications
@@
We first fit an implicit SDF volume to multi-view renders of the mesh, then apply our gradient with the same schedule as in text-based NeRF optimization. We lower the learning rate for geometry parameters to prevent divergence from the guiding sketch. Holding other hyperparameters equal, we compare our gradient and the SDS gradient in Figure A1.
@@
Figure A2: Diffusion illusions. We generate overlaid optic illusions with SDS and our method. While SDS suffers from color artifacts, our methods produce more details and proper color.
@@
Prior works have shown that diffusion models can be leveraged to generate optical illusions [17, 5]. In these settings, the same image looks semantically different when transformed. To use the diffusion model sampling process, a previous study shows that the transformation has to be orthogonal [17]. However, there remain interesting illusions that are not formed by orthogonal transformation. One such is the rotation overlays. Given a base and a rotator image, by composing the base image with the rotator image at different angles, rotation overlays use two images to display four images. As such composition is not defined by an orthogonal matrix, the existing method [5] employs SDS to optimize the base and rotator images. Such a method suffers from the over-saturation problem, as shown in Figure A2. We show that our method can generate such optical illusions with better visual quality.
(a) MVDream [57]
(b) SDXL [47]
Figure A3: Comparison of SDS and ours with MVDream [57] and SDXL [47]. We compare SDS with our two-stage process in two new settings (MVDream and SDXL). The two-stage process produces more natural colors and realistic details.
B.2 Additional Qualitative Results
Additional text-to-image results.
We explore our proposed method across different base models in text-to-image experiments, including MVDream [57] and SDXL [47]. Since MVDream denoises four camera-conditioned images jointly, we treat the canvas of four images as a single optimization variable for the SDS gradient. In Figure A3, we compare the SDS baseline to the proposed two-stage optimization, in which we generate more natural colors and detail. This is especially noticeable in the background around the crocodile and donkey.
Figure A4: Comparison with more text-to-3D baselines. We apply our two-stage optimization as a drop-in replacement of SDS in Fantasia3D [9], Magic3D [35] and CSD [76] for texture refinement. We notice that this change greatly improves details and visual quality and reduces SDS artifacts.
@@ B.2 Additional Qualitative Results
For text-guided NeRF optimization comparison against baselines, we show more results in Fig. A7. We test on the prompts used in the original paper [70] and additional prompts [71] that we find to be challenging. We notice that SDS often suffers from over-saturation problems. Our method does not require training a LoRA while it can still improve SDS by getting rid of the color artifacts and generating more details.
We also perform comparisons with more competitive baselines. We test with Fantasia3D [9], Magic3D [35], and CSD [76] through a drop-in replacement of SDS with our method. Specifically, all three methods optimize a textured DMTet, which is initialized from an SDS-optimized NeRF, using SDS or CSD for 5k or 10k iterations. We replace the SDS or CSD stage of these approaches with the two-stage optimization motivated by our framework. Just like our text-to-3D NeRF experiment, we perform the first stage for 60% of iterations and the second stage for 40% of iterations. Note that we keep all the other hyperparameters the same, which were tuned for the baselines, not our method. This replacement leads to the same optimization time as the original methods. For Fantaisia3D and Magic3D, we use threestudio for fair comparison (Magic3D does not have code available) and the default prompts, which are generally believed to work the best with this reimplementation. For CSD, we use the official implementation. As shown in Figure A4, our method improves the visual quality of all the methods by reducing the oversaturated artifacts of SDS and improving the details.
Figure A5: Ablation study of negative prompts. We compare SDS results with those from the two-stage optimization (bridge) using our negative prompts and five sets of negative prompts generated by GPT (GPT 1 is the first set, GPT 2 second, GPT 3 third, etc.). All negative prompts produce similar results and outperform the SDS baseline.
B.3 Ablation Study
Ablation study of the negative prompts.
We explore how the choice of negative prompts in our proposed methods affects the optimization. We prompted GPT-4 through ChatGPT a single time to generate alternative negative prompts using the following:
Here’s a set of "negative prompts" to append to a text-to-image prompt that describe undesirable image characteristics: ", oversaturated, smooth, pixelated, cartoon, foggy, hazy, blurry, bad structure, noisy, malformed" I want to try a variety of them, please brainstorm many of roughly the same length.
We produce five variants through these methods as the alternative negative prompts:
  1. 1.
    ", washed out, grainy, distorted, flat, smeared, overexposed, undefined, choppy, glitchy, dull"
  2. 2.
    ", low contrast, jumbled, faint, abstract, over-sharpened, muddy, cluttered, vague, jagged, poor detail"
  3. 3.
    ", soft focus, muffled, streaky, patchy, ghosted, murky, unbalanced, skewed, mismatched, overcrowded"
  4. 4.
    ", overbright, scrambled, bleary, blocky, misshapen, uneven, fragmented, obscured, chaotic, messy"
  5. 5.
    ", dull tones, compressed, smeary, out of focus, unrefined, lopsided, erratic, irregular, spotty, stark"
We keep other hyperparameters identical and only ablate the negative prompts with the variations. As shown in Figure A5, we do not see obvious differences between our prompts and the variants.
Figure A6: Ablation study of our method without stage 1. We show directly optimizing with $y_{\text{src}}$ from the start can undermine the quality of the geometry and produce unnecessary content.
@@ B.3 Ablation Study
Instead of switching to stage 2 during the optimization process, we ablate with starting without any SDS optimization from the beginning. That is, we always use the $y_{\text{src}}$ with the descriptors “, oversaturated, smooth, pixelated, cartoon, foggy, hazy, blurry, bad structure, noisy, malformed”. As shown in Figure A6, this makes it hard to generate the proper geometry even though the local texture looks reasonable and is inclined to produce excessive details that are not described by the texts. We suspect that this is because using $y_{\text{src}}$ increases the mismatching error at the beginning of the optimization process when the initialization does not resemble the target prompt at all.
Figure A7: Additional comparison of text-guided NeRF optimization. We show more examples to compare with different distillation methods, SDS and VSD.
Appendix C Potential Social Impact